What is the probability that at least 2 people in a group of 8 people have the same birthdate? (Assume every year has exactly 365 days - i.e., there is no Leap Day.)

Answer by ikleyn(52798)

.
What is the probability that at least 2 people in a group of 8 people have the same birthdate?
(Assume every year has exactly 365 days - i.e., there is no Leap Day.)
~~~~~~~~~~~~~


This probability is the COMPLEMENT to the probability that ALL 8 people have DIFFERENT birthdays.


The probability that 8 people have different birthdays is


    Q =  = 


             //  (the product of 7 (seven)  multipliers) = 0.925665...


Thus the probability under the problem's question is


    P = 1 - Q = 1 -  = 1 - 0.925665... = 0.0743  (rounded).    ANSWER

Solved.

----------------

See the lesson
    - Coinciding birthdays
in this site.